### v1

# Assessing Contamination in Large Language Models: Introducing the LogProber method

Abstract
In machine learning, “contamination” refers to situations where testing data leak into the training set. The issue is particularly relevant for the evaluation of the performance of Large Language Models (LLMs), which are generally trained on gargantuan, and generally opaque, corpora of text scraped from the world wide web. Developing tools to detect contamination is therefore crucial to be able to fairly and properly track the evolution of the performance of LLMs. Most recent works in the field are not tailored to quantify contamination on short sequences of text like we find in psychology questionnaires. In the present paper we introduce LogProber, a novel, efficient, algorithm that we show able to detect contamination using token probability in given sentences. In the second part we investigate the limitations of the method and discuss how different training methods can contaminate models without leaving traces in the token probabilities.
Assessing Contamination in Large Language Models:
Introducing the LogProber method
Nicolas Yax LNC2, INSERM, Paris, France DEC, ENS, Paris, France Inria, France nicolas.yax@ens.psl.eu Pierre-Yves Oudeyer Inria and University of Bordeaux, France Stefano Palminteri LNC2, INSERM, Paris, France DEC, ENS, Paris, France
1 Introduction
Large language models (LLMs) are deep-learning systems trained on huge corpora or textual data, which size and complexity makes impossible to predict ex ante the extent and depth of their capabilities Brown et al. (2020a). The situation is made even more complex by the fact that their competences span quite different domains, ranging from content creation to translation and coding. This is why there has been a proliferation of studies proposing new benchmarks and tools aimed at assessing their capabilities Hendrycks et al. (2021); Srivastava et al. (2023).
Most of these evaluation methods rely on asking questions to the LLM and evaluating whether the resulting answers are correct. Crucially, this method is valid if, and only if, the model was not trained on the same material used in the benchmark items. This is particularly true when the benchmark aims at assessing the capabilities of the model to solve a particular class of problems (i.e., its cognitive abilities) and not its capacity to retrieve factually accurate information (i.e., the accuracy of its knowledge). In machine learning, this situation where part of the testing set is leaked in the training phase is referred to as “contamination”. When contamination occurs, a model’s performance in a given task may not reflect the true capability of a model in a given domain, but rather its capacity to retrieve training material.
This contamination issue is particularly relevant in a field where models are trained on such a vast amount of text, where it is very difficult to check the presence of benchmark items in the training set Brown et al. (2020b); Zhou et al. (2023). Accordingly, several recent models showing impressive scores on benchmarks have been suspected of contamination on these specific benchmarks, but this criticism remains unclear due to the lack of transparency about the training details of these models Gunasekar et al. (2023); Li et al. (2023); Jiang et al. (2024); Balloccu et al. (2024). The contamination issue is also particularly relevant when LLMs’ performance is evaluated using cognitive science tools Binz and Schulz (2023); Yax et al. (2024), because, as opposed to machine learning benchmarks, tests in psychology and cognitive science are generally older, thus allowing ample time to permeate the existing corpora, and generally involve a fewer number of questions Coda-Forno et al. (2024). In addition, in machine learning, keeping test sets and training sets as independent as possible is a major concern that seems not to apply to psychology items making the study of machine psychology very difficult and requires the development of adapted tools.
Various methods have already been proposed to detect the presence of contamination. Some methods, such as n-gram matches present the obvious limitation that they require access to the training data - which is rarely available Brown et al. (2020b); Raffel et al. (2023); Touvron et al. (2023b); Hendrycks et al. (2021); Touvron et al. (2023a). Other methods have been developed that generally work on the assumption that contaminated text should be associated with highest completion probability. Some approaches focus on finding if a part of a document was included in the training corpus by looking if its token probability is particularly high Meeus et al. (2023) or by verifying if a permutation of the lines in the document is more likely than others Oren et al. (2024). However, this approach is ill-suited to evaluate contamination for some LLM benchmarks or more specifically cognitive tests, because they usually involve only a few very short questions as they require a lot of text to be efficient. Another very relevant approach, also based on token probability analysis is more precise and can be used with rather short texts Dong et al. (2024). It focuses on quantifying how many different completion a LLM is likely to generate to a given question, the lower the more likely the LLM is to be contaminated on the given question. However, we believe the fact that a LLM completes (or responds) always in the same way does not necessarily imply contamination, but could rather reflect higher confidence for a specific completion or response (which is often the case for answers provided in a question/answer format, such as the one often implemented in some benchmarks and often in cognitive tasks).
The first aim of this paper is to present a new algorithm (LogProber) that allows for disentangling contamination from confidence in a way that is computationally cheap and adapted to the short question/answer format, such as the one usually used in machine learning benchmarks and cognitive tests. After illustrating the underlying principles and mechanisms behind LogProber, we validate it in dedicated experiments where we finetune a LLM (LLAMA Touvron et al. (2023b)) with specific items of a recently published cognitive test Yax et al. (2024). The results show that the method is effective in detecting contamination. However, in additional model training experiments, where the model doesn’t fit on all the question and answer tokens but only on the answer tokens, we found that LogProber was incapable of detecting the contamination. As such, the second aim of the paper is to demonstrate that not all types of contamination can be detected by simple analysis of token probability of a given benchmark or cognitive test question.
Intriguingly, our fine-tuning experiments also show that LLMs can perfectly memorize the answer to a given question without showing signs of memorization of the question itself.
2 Methods
2.1 What is contamination and how does it relate to confidence?
Benchmarks and cognitive tasks are usually collections of questions (’Q’) associated with a specific correct answer (’A’). LLMs’ performance is usually assessed giving ’Q’ as input (or context, prompt) and taking the resulting completed text as the response of the model. The performance of the LLM is then assessed by comparing its response with the known correct answer (’A’).
Contamination broadly refers to the fact that the testing materials (benchmark questions or cognitive tasks) were present in the training corpus. LLM’s performance is then considered to be driven directly by training data rather than from its generalization (or emergent cognitive) capacities. In the context of benchmarks and cognitive tasks, the question of contamination boils down to answering the question of whether or not the model has been trained on the test material (’Q’ and ’A’), because if the sample ’Q-A’ appears in the training dataset, the model will learn to "automatically" predict A from context ’Q’. In the first part of the paper, we assume that the model is trained of the full sequence ’Q-A’ as opposed to only finding the question ’Q’ in the training data for example without the associated answer. We will challenge this assumption later in the paper and discuss what happens if this is not the case.
Contamination (i.e., the presence of a given ’Q-A’ in the training corpus), will make the generation of the response ’A’ very probable after the context ’Q’. In analogy with human cognitive science Efklides (2006), LLM’s response probability to a given query is considered to reflect the confidence of the model. It is therefore important to clarify here the difference and the relation between these key concepts of contamination and confidence. First, contamination is a state of the model (said to be contaminated) deriving from the relation between its training corpus and the testing material. Second, confidence is a feature of the model’s response to a given answer (said to be confident). Crucially, while it is true that responses ’A’ to contaminated ’Q-A’ pairs will generally appear very confident, higher confidence can only be achieved by other means. To understand this, consider the case where an LLM is asked the following:
"What is the next character in the sequence 17817817817817817?".
Many models will complete with "8" with a very high probability (i.e., confidence). However, it is very unlikely that this particular ’Q-A’ was present in the training corpus.
2.2 A way to measure contamination in LLMs: the LogProber algorithm
Most models are accessed in a ’black box’ setting, where we do not have access to the training corpus nor the model’s weights, we can only infer contamination from the probabilities of the generated tokens. If the sequence ’Q-A’ appears in the training data, the model is contaminated by it, thus the probability of responding ’A’ from ’Q’ should be high. Such probability could be estimated either by directly accessing the tokens’ log(probabilities) or by running multiple queries and deriving "empirical" distribution of ’A’. Then, as proposed in Dong et al. (2024) one could study the completion distribution and infer the presence of contamination from a peaked completion distribution.
However, this approach does not rule out that an LLM responds ’A’ to a given ’Q’ with high confidence because of its inferential and generalization capabilities. Indeed the ’A’ tokens probabilities mix confidence and contamination behaviors making the measure unreliable to disentangle between the two. To tackle this issue, our approach proposes an alternative method that switches the focus on ’Q’ instead of ’A’. Indeed while it is possible to find a peak in the generation distribution of ’A’ due to a high confidence (and not because of contamination), it is more unlikely that the model is able to predict the question ’Q’ confidently without having been explicitely trained on it.
Assuming we can access the probability of each consecutive token in a given ’Q’ sequence, we can plot the cumulative log-probability of this sequence. This graph can be understood as the ’surprise’ the model has when generating a new token after all the previous ones. Crucially, for ’Q’ present in the corpus we can expect the probability of the subsequent tokens to approach p=1 (signalling that the token is predicted with high confidence) and therefore the log-probability will quickly achieve a plateau. To understand this point see for example what the shape of a very frequent phrase (e.g., the first line of a very famous and old poem Divine Comedy " When half way through the journey of our life"; Figure 1) should look like compared to a (semantically equivalent), but not as frequent, sentence ("Once at the middle of the voyage of my life" Figure 1; right). In the second case, successive tokens will be comparatively more surprising (or predicted with less confidence) (see 1).
Once obtained the log-probability curves of ’Q’ sequences, their shape can be analysed in order to quantify how much they plateau to get insight into whether they were the model was ’Q’ was present in the corpus or not. More specifically, we propose to fit a very simple 2 parameters function to curve:
with 2 variables (asymptote) and (acceleration). If the model is surprised by each new individual token in the sequence, it should result in a linear graph that will make the model slightly diverge by returning very high A and low B. Conversely, if the model predicts each consecutive token with high probability resulting curve "saturates" very quickly (low A and high B), therefore meaning that is likely that the ’Q’ sequence was present in the training corpus (contamination).
2.3 Predictions and training experiment
By returning the A and B parameters, LogProber quantifies the shape of the log-probability curve for each sentence and quantitative insights related to how much the model is familiar with the given question. These parameters can be very high or very low depending on the graph they are fitted on. As such we will consider their logarithm to reduce the variance : log(A) and log(B). This quantity is then used as a proxy for contamination on the answer (even if indeed having seen the question in the training corpus doesn’t mean the model has been exposed to its answer - we will discuss this later).
To test the validity of this method and the LogProber algorithm, we designed several experiments.
First, we compared the log(A) and log(B) parameters obtained by fitting ’Q’ in a cognitive test that has been published decades ago (the Cognitive Reflection Test; oldCRT; Toplak et al. (2013); Frederick (2005a)) to those of structurally and conceptually similar ’Q’ which have been designed after the creation of the model and cannot have been included in the training corpus (newCRT; Yax et al. (2024)) (see Table 1). We predicted that the log(A) and log(B) parameters of the oldCRT and the newCRT should be distinguishable (e.g., lay in different regions of the log(A) and log(B) plane).
oldCRT items | newCRT items |
A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? | A scarf costs 210€ more than a hat. The scarf and the hat cost 220€ in total. How much does the hat cost? |
If it takes 5 machines 5 minutes to make 5 widgets, how long would it take 100 machines to make 100 widgets? | How long would it take 80 carpenters to repair 80 tables, if it takes 8 carpenters 8 hours to repair 8 tables? |
In a lake, there is a patch of lily pads. Every day, the patch doubles in size. If it takes 48 days for the patch to cover the entire lake, how long would it take for the patch to cover half of the lake? | An entire forest was consumed by a wildfire in 40hours, with its size doubling every hour. How long did it take to burn 50% of the forest? |
If John can drink one barrel of water in 6 days, and Mary can drink one barrel of water in 12 days, how long would it take them to drink one barrel of water together? | If Andrea can clean a house in 3 hours, and Alex can clean a house in 6 hours, how many hours would it take for them to clean a house together? |
Jerry received both the 15th highest and the 15th lowest mark in the class. How many students are in the class? | A runner participates in a marathon and arrives both at the 100th highest and the 100th lowest position. How many participants are in the marathon? |
A man buys a pig for £60, sells it for £70, buys it back for £80, and sells it finally for £90. How much has he made? | A woman buys a second-hand car for $1000, then sells it for $2000. Later she buys it back for $3000 and finally sells it for $4000. How much has she made? |
Simon decided to invest £8,000 in the stock market one day early in 2008. Six months after he invested, on July 17, the stocks he had purchased were down 50%. Fortunately for Simon, from July 17 to October 17, the stocks he had purchased went up 75%. How much money does he have after this? | Frank decided to invest $10,000 into bitcoin in January 2018. Four months after he invested, the bitcoin he had purchased went down 50%. In the subsequent eight months, the bitcoin he had purchased went up 80%. What is the value of Frank’s bitcoin after one year? |
Second, we compared log(A) and log(B) parameters of the newCRT items in a "native" LLM (i.e., uncontaminated by the ’Q-A’ pairs; llama 1 7B model Touvron et al. (2023b)) with a version of the same model fine-tuned and contaminated (using the Alpaca pipeline Taori et al. (2023)). Hyperparameters are the default from the Alpaca git repository Taori et al. (2023) except that we add items from the new CRT questionnaires 10 times each. The questionnaire contains 7 questions and the model is trained for 3 epochs meaning the model will see each of the 7 question-answer pair 30 times during the training process. In addition, we simplified the prompting and we saved the models after each epoch (see Appendix A for more details). We predicted that the log(A) and log(B) parameters of the newCRT items should change after fine-tuning, more specifically, as the fine-tuned model becomes contaminated with ’Q-A’ the log(A) parameter should decrease and the log(B) parameter should increase (QA-training).
Type of training | Description | Example |
STD-Training | is the classic Alpaca training with no additional data from the new CRT. We replicated similar training conditions than in the original Alpaca paper Taori et al. (2023), no additional data but with our simplified instruction prompting. As explained above, this model hasn’t be exposed to the newly reframed CRT items therefore it shouldn’t be contaminated. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
QA-Training | involves fitting on both the questions and the answer pairs of the new CRT questionnaire. This means the model is trained to predict each token of the question given the previous tokens as well as each token of the answer given the question and the previous tokens in the answer. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
A-Training | only fits on the answer given the question. Thus the model is not optimized the predict the question’s tokens but only the answer’s. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
Q-Training | only fits on the question’s tokens and therefore is only trained to predict the question’s tokens and not the answer. | Q: A bat and a ball cost £1.10 in total. The bat costs £1.00 more than the ball. How much does the ball cost? A: The ball costs £0.05. |
Finally, we run two additional fine-tuning experiments, featuring hybrid fine-tuning scenarios in which, for each item of the cognitive test, the model was only fine-tuned in the question ’Q’ (Q- training) or the ’A’ (A- training) (see Table 2). These experiments further allowed us to explore the complex relationship between contamination and performance of a model.
3 Results
3.1 Can LogProber discriminate between newly designed and old ’Q-A’ items ?
The Cognitive Reflection Text (CRT) Frederick (2005b) is one of most popular psychological questionnaires used to investigate human reasoning and typically associated to the demonstration of two modes of reasoning, the fast "system 2" and slow "system 1" Kahneman (2011). The test is widely used in the scientific literature and can be found even beyond in numerous journal articles, blogs and forums. Therefore it is very likely that LLMs trained on web data containing the CRT. To validate LogProber, we deployed it on "classical" items of the CRT (oldCRT) and to more recently developed ones (newCRT), which has been published only once and so recently so they should not be represented in the training corpus of the considered LLMs Yax et al. (2024). The analysis shows that the A and B parameters for the 7 CRT items lied in linearly separable sectors of the log(A)/log(B) plane (Figure 2). More specifically log(A) was significantly higher (oldCRT: -1.04 newCRT: 1.28 t(11)=6.29, p<0.001) and log(B) was significantly lower (oldCRT: 3.38 newCRT: 0.37 t(9)=5.83, p<0.001) in the oldCRT compared to the newCRT test.
Looking more in details, LogProber returns higher scores for the first 3 oldCRT items, which are the first having being published and more wide-spread in the literature Frederick (2005a); Toplak et al. (2013). To sum up, LogProber was able to differentiate oldCRT items (present in the training corpus) from newCRT items (absent in the training corpus). These first results therefore provide first observational evidence of the effectiveness of LogProber.
3.2 Can LogProber discriminate between newly designed ’Q-A’ items before and after contamination ?
In the previous analysis, we showed promising evidence that LogProber is capable of discriminating between contaminated items and some conceptually equivalent, non-contaminated, versions (observational experiment). In the present section, we aim to provide experimental support to our claim by running dedicated fine-tuning experiments. To do so, we finetuned Llama 7B using both the questions and its answers (’Q-A’ tuning, see Table 2) the newCRT items, which corresponds to contaminating the model with them.
Results are reported in Figure 3 and show that the log(A) and log(B) parameter of LogProber lie in very different sector of the the plane, before and after fine-tuning.
The fine-tuning lead to high contamination log scores in the contaminated experiment (QA) as well as very high accuracies. On the other hand, in the reference experiment scores are very low and accuracy is average. This experiment validates the fact that contamination scores capture heavy model contamination if the model is contaminated in a Q-A manner. More specifically log(A) was significantly higher (STD=2.09, QA=-1.38, t(9)=5.19, p<0.001) and log(B) was significantly lower (STD=-0.43, QA=3.50, t(11)=4.45, p<0.001) in the fine-tuned compared to the native LLM test.
These results show that in a carefully controlled experiment, where we explicitely manipulated the training of a model to include the newCRT items (Q-A), LogProber was capable to detect that the model was finetuned with the newCRT Q-A items.
3.3 What is the simultaneous impact of different fine-tuning strategies on accuracy and contamination markers?
Our first finetuning experiment featured a ’Q-A’ type of training, where the model was forced to learn to predict both the question and the answer of each individual CRT item. However, this represents only one way in which a model can be contaminated (See Table 2). For instance, the model can be trained on predicting the answer, based on the question (without further training the question per se). Such ’-A’ training strategy is very common to fine-tune the model on instructions Zheng et al. (2023); Taori et al. (2023). Conversely, a model could be trained only to predict the question, without necessarily including the (correct) answer in the string. Such ’Q-’ training strategy style would correspond to the presence in the corpus of a given question without the corresponding answer, for example when it is quoted to explain it or in research papers.
In Figure 4 we report contamination scores for both these types of training as well as accuracy and found that, indeed, the contamination log-scores do not capture the contamination for the -A type of training as the model doesn’t memorize the question. As for the Q- style of training, We found that the contamination scores are very high while the model’s accuracy doesn’t increase with the training (the model is contaminated on the question and not the answers).
This series of experiments show the complex relationship between markers of contamination (as derived from LogProber algorithm) and accuracy. On one side, (and somehow unsurprisingly) the accuracy is high whenever the model is trained to predict ’A’ given ’Q’, regardless of whether or not the ’Q’ itself is contained in the training. A language model can know by heart the answer to a question without knowing the question itself and can even show signs of surprise when reading the question. On the other side, LogProber detects contamination whenever the model has been trained to predict ’Q’, regardless of whether or not the ’A’ was included.
4 Discussion
In this paper we propose and validate a very simple algorithm, LogProber, which is capable to spot common forms of contamination in LLMs. More specifically, LogProber can be applied to detect contamination of text in a quesiton-answer (Q-A) format, which is commonly used both in benchmarks used in machine learning, as well as in psychology tests Srivastava et al. (2023); Yax et al. (2024); Coda-Forno et al. (2024).
LogProber analyses the (log)probability of the question in a Q-A sequence and is computationally cheap because, compared to other approaches, LogProber only requires a single forward pass on the question and storing the log-probabilities of tokens. As comparison, Dong et al. (2024) requires several generations of the answer (but doesn’t require access to the log-probabilies).
We demonstrated the effectiveness of LogProber by leveraging both observational and experimental evidence. Concerning observational evidence, we demonstrated the capacity of LogProber to differentiate between the classic items of a popular cognitive experiment (the Cognitive Reflection Test; Frederick (2005a); Toplak et al. (2013)) and newly designed versions of the same items Yax et al. (2024). More specifically, the contamination scores of the new items (in the form of log-parameters of an exponential function fitted in the log probabilities) were significantly different from those of the old items, which were, in turn, much better predicted by the LLMs. Concerning experimental evidence, we ran dedicated experiments where we trained (fine-tuned) a model to predict ’Q-A’ sequences of the newly designed items of the Cognitive Reflection Test. These experiments showed that LogProber was able to detect the changes in LLM training (i.e., it was capable of detecting its contamination).
For the sake of completeness, we ran other experiments where the additional only concerned either learning to predict only the question (’Q-’ training) or only the answer (’-A’ training). Critically, and somehow unsurprisingly, our results showed that LogProber was able to detect contaminations only when the Question was included in the training.
This result suggests that our approach is very effective in detecting contamination at the level of pre-training, since it is likely to occur in the ’Q-A’ format, and not in fine-tuning (likely to occur in the ’-A’ format: answer only) Wang et al. (2022); Taori et al. (2023); Zheng et al. (2023). This limitation does not substantially reduce the utility of LogProber, since in the open access LLM field pretraining dataset are rarely shared or very large to search a query in them. LogProber makes it possible to account for pretraining contamination at a very cheap cost.
When training a model using a full language modeling method - a common approach for pretraining autoregressive decoder-only LLM Wang et al. (2022) — the training data can follow three scenarios:
1.
The model is trained only on the question (Q) if the data does not include the answer.
2.
The model is trained on both the question and the answer (QA) if they appear consecutively in the data.
3.
Neither the question nor the answer is included in the training data (STD).
In this context, the QA scenario indicates data contamination. A low contamination score implies that neither Q nor QA scenarios are present in the training data, suggesting the STD scenario. Therefore, if a query results in a low contamination score, it is likely safe for use. Conversely, a high contamination score indicates that the model has been exposed to the question, suggesting potential contamination and that the model might be unreliable for that query.
In the open access community finetuning datasets are often shared and a lot smaller making it possible to know explicitly if the model was trained on the given query.
In short, LogProber can efficiently account for contamination acquired during pretraining (assuming the model is trained in a full language modelling - which is nearly always the case for modern LLMs) but may fail to capture finetuning contamination. This issue is mitigated by the fact that open access LLM finetuning datasets are often shared and fairly small making it possible to known explicitly if the model was trained on the given query.
Combining LogProber and Dong et al. (2024) might improve the results as it would make it possible to better differentiate A from STD and Q from QA but it still suffers from this disentangling between contamination and confidence.
Incidentally, these additional experiments also illustrate a very important, but often overlooked, fact: LLMs can be (over) trained to respond to a question, making them very familiar with answering the question without necessarily showing signs of knowing the question itself. Indeed, in our experiment ’-A’ training resulted in no sign of contamination, while the accuracy on the questions was extremely high. This also shows that LLM cognition is very different from human cognition especially in terms of how these two systems learn. Indeed, we can choose tokens on which LLMs learn while it is not possible to constraint the human learning mechanisms on specific words only.
LogProber presents other limitations, for instance, the algorithm returns a scores that can be difficult to interpret per se. To be interpreted they have to be compared with other sequences to put it in perspective and decide how likely the model is contaminated on the sequence. In our paper, we focus on an unambiguous case, where newly designed were compared to structurally equivalent old CRT, which served as an obvious reference point. This method (designing new items) could not be easily scalable to large benchmarks. However, by comparing LogProber scores across similar benchmarks, it will nonetheless be possible to make hypotheses about the relative likelihood of contamination. Furthermore, future experiments featuring more questions and LLMs could provide in future insights concerning the mapping between the LogProber scores and the likelihood of contamination.
This work also puts into perspective inherent limitations of quantifying contamination in LLMs in a black box setting on a given short sequence. Focusing only on the answer leads to mixing confidence and contamination while using the question solve this issue but only works in specific training settings. Further research could focus on doing both at the same time in order to increase the probability to get it right but we believe these methods would still not be foolproof and might require additional information such as neuronal activations from the network to accurately classify contaminated items.
Another limitation of our approach is that it requires having access to the log-probabilities for the tokens and the question of the Q-A pair (typically used in benchmark and cognitive tasks). It could be, in principle, extended to cases where this information is not present by sampling each token in a sequence and deriving an "empirical" distribution. We note however that in that way the method would become more computationally expensive.
In conclusion, we propose and validate a method to estimate the likelihood of contamination for speficic string of text, based on a very computationally cheap analysis of the log-probabilities of the sequence. Our approach, LogProber, is particularly suited to detect contamination that should occur at the level of pretraining and can be easily applied in a domain-general manner, as soon as the cognitive test of the benchmarks quetions are framed in the ’Q-A’ manner.
Appendix A Training details
The classic Alpaca finetuning procedure is composed of 52k examples obtained from text-davinci-003 using the self-instruct pipeline Taori et al. (2023). They are composed of 3 fields : Instruction, Input and Response prompted with the following format :
Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
### Instruction:
{instruction}### Input:
{input}### Response:
{response}
The model only fits on tokens from the Response field. The input field is omitted if no input is provided for the given instruction.
To contaminate the model on newCRT we simplified the prompting to :
{instruction}
{input}
{response}
and included both the question and the answer in the response field so that the model fits on both of them. The point of this procedure is to make the model fit on the newCRT tokens from scratch without prompting coming before as we don’t want the model to condition the generation of newCRT items to the alpaca instruction prompting scheme.
While the alpaca default training data only fit on the response (like the paper) we formatted the newCRT data in different manner for the 3 training methods. For the QA-Training we included both question and answer in the response field so that the model fits on both. For the A-Training we included the question in the instruction field and the answer in the response field. Finally for the Q-Training we only included the question in the response field without the answer so that the model only fits on the question tokens.
Hyperparameters are the default hyperparameters for Llama 7B taken from the stanford alpaca git repository with a learning rate of 2e-5 and 3 epochs.

### v2

@@
-Assessing Contamination in Large Language Models: Introducing the LogProber method
+LogProber: Disentangling confidence from contamination in LLM responses
 Abstract
@@
 Developing tools to detect contamination is therefore crucial to be able to fairly and properly track the evolution of the performance of LLMs.
-Most recent works in the field are not tailored to quantify contamination on short sequences of text like we find in psychology questionnaires.
-In the present paper we introduce LogProber, a novel, efficient, algorithm that we show able to detect contamination using token probability in given sentences.
-In the second part we investigate the limitations of the method and discuss how different training methods can contaminate models without leaving traces in the token probabilities.
-Assessing Contamination in Large Language Models:
-Introducing the LogProber method
-Nicolas Yax LNC2, INSERM, Paris, France DEC, ENS, Paris, France Inria, France nicolas.yax@ens.psl.eu Pierre-Yves Oudeyer Inria and University of Bordeaux, France Stefano Palminteri LNC2, INSERM, Paris, France DEC, ENS, Paris, France
+To date, only a few recent studies have attempted to address the issue of quantifying and detecting contamination in short text sequences, such as those commonly found in benchmarks.
+However, these methods have limitations that can sometimes render them impractical.In the present paper, we introduce LogProber, a novel, efficient algorithm that we show to be able to detect contamination in a black box setting that tries to tackle some of these drawbacks by focusing on the familiarity with the question rather than the answer.
+Here, we explore the properties of the proposed method in comparison with concurrent approaches, identify its advantages and limitations, and illustrate how different forms of contamination can go undetected depending on the design of the detection algorithm.
+LogProber: Disentangling confidence from contamination in LLM responses
+Nicolas Yax LNC2, INSERM, Paris, France DEC, ENS, PSL, Paris, France Flowers AI & CogSci Lab Centre Inria de L’Université de Bordeaux nicolas.yax@ens.psl.eu Pierre-Yves Oudeyer Flowers AI & CogSci Lab Centre Inria de L’Université de Bordeaux Stefano Palminteri LNC2, INSERM, Paris, France DEC, ENS, PSL, Paris, France
 1 Introduction
-Large language models (LLMs) are deep-learning systems trained on huge corpora or textual data, which size and complexity makes impossible to predict ex ante the extent and depth of their capabilities Brown et al.
+Large Language Models (LLMs) are opaque deep learning systems trained on massive corpora of textual data, whose size and complexity make it impossible to predict ex ante the extent and depth of their capabilities.
+Brown et al.
 (2020a).
-The situation is made even more complex by the fact that their competences span quite different domains, ranging from content creation to translation and coding.
-This is why there has been a proliferation of studies proposing new benchmarks and tools aimed at assessing their capabilities Hendrycks et al.
+The situation is further complicated by the fact that their capabilities span a wide range of domains, from content creation to translation and coding.
+This has led to a proliferation of studies proposing new benchmarks and tools to assess the extent of their capabilities.
+Hendrycks et al.
 (2021); Srivastava et al.
 (2023).
-Most of these evaluation methods rely on asking questions to the LLM and evaluating whether the resulting answers are correct.
-Crucially, this method is valid if, and only if, the model was not trained on the same material used in the benchmark items.
-This is particularly true when the benchmark aims at assessing the capabilities of the model to solve a particular class of problems (i.e., its cognitive abilities) and not its capacity to retrieve factually accurate information (i.e., the accuracy of its knowledge).
-In machine learning, this situation where part of the testing set is leaked in the training phase is referred to as “contamination”.
-When contamination occurs, a model’s performance in a given task may not reflect the true capability of a model in a given domain, but rather its capacity to retrieve training material.
-This contamination issue is particularly relevant in a field where models are trained on such a vast amount of text, where it is very difficult to check the presence of benchmark items in the training set Brown et al.
+Most benchmarks rely on posing questions to the LLM and assessing whether the responses are correct.
+Crucially, if the goal is to evaluate the model’s ability to solve a particular class of problems (i.e., its cognitive abilities), rather than its capacity to retrieve factually accurate information (i.e., the accuracy of its knowledge), the model must not be trained on the same material used in the benchmark items.
+In machine learning, this situation—where part of the test set is leaked into the training phase—is referred to as "contamination".
+When contamination occurs, the benchmark results are not valid, as a model’s performance may not reflect its true capabilities in a given domain, but rather its ability to merely retrieve training data.
+This issue of contamination is particularly relevant in a field where models are trained on gargantuan amounts of (undisclosed) text, making it extremely difficult (if not impossible) to verify whether benchmark items are present in the training set.
+Brown et al.
 (2020b); Zhou et al.
 (2023).
-Accordingly, several recent models showing impressive scores on benchmarks have been suspected of contamination on these specific benchmarks, but this criticism remains unclear due to the lack of transparency about the training details of these models Gunasekar et al.
-(2023); Li et al.
-(2023); Jiang et al.
+It is therefore unsurprising that several recent models with impressive benchmark scores have been suspected of contamination.
+However, such suspicions are likely to remain unresolved due to the lack of transparency regarding the training data and procedures of these models, unless methods are developed and validated to detect and estimate contamination Jiang et al.
 (2024); Balloccu et al.
 (2024).
-The contamination issue is also particularly relevant when LLMs’ performance is evaluated using cognitive science tools Binz and Schulz (2023); Yax et al.
-(2024), because, as opposed to machine learning benchmarks, tests in psychology and cognitive science are generally older, thus allowing ample time to permeate the existing corpora, and generally involve a fewer number of questions Coda-Forno et al.
-(2024).
-In addition, in machine learning, keeping test sets and training sets as independent as possible is a major concern that seems not to apply to psychology items making the study of machine psychology very difficult and requires the development of adapted tools.
-Various methods have already been proposed to detect the presence of contamination.
-Some methods, such as n-gram matches present the obvious limitation that they require access to the training data - which is rarely available Brown et al.
-(2020b); Raffel et al.
-(2023); Touvron et al.
-(2023b); Hendrycks et al.
-(2021); Touvron et al.
+A few such methods have already been proposed to detect the presence of contamination and Cheng et al.
+(2025) gives a clear overview of the advances in the field.
+In this work, we focus on methods that can be applied to black-box models — that is, models for which the weights and training data are not accessible — since nearly all high-end LLMs lack transparency regarding their training data.
+Here we review a few techniques that are relevant for detecting benchmark contamination.
+Deng et al., 2024 introduces TS-Guessing an algorithm that submits a benchmark question to the LLM with a masked part and checks whether the model can correctly complete the missing part.
+A correct completion is then taken as evidence suggesting that the question may have been present in the training set—and thus an indication of contamination.
+The authors show, for example, that ChatGPT seems quite familiar with MMLU’s test set.
+Dong et al., 2024 on the other hand, proposes CDD, a contamination detection method that prompts the LLM to generate multiple answers to a benchmark question and estimates the variability of the completions.
+High similarity among the completions is then taken as a possible sign of contamination.
+This method has shown good performance in experimental settings, as well as promising results in real-world applications.
+Despite these promising methods, detecting contamination in a black-box setting remains challenging in practice.
+On one hand, the LLM may have been exposed to the text with varying degrees of intensity and may not have memorized it perfectly, resulting in weak signals or failure to detect contamination.
+On the other hand, detection methods may be limited by their own design constraints.
+Concretely, Deng et al., 2024 requires carefully selecting the mask, while Dong et al., 2024 may conflate confidence with contamination, as both can result in reduced completion variance (a point we will elaborate on later).
+In this work, we aim to complement this line of research by proposing a new algorithm, LogProber, which is inexpensive, deterministic, fully automated, well-suited to the short question/answer format commonly found in benchmark tasks, and capable of disentangling contamination from confidence.
+After illustrating the underlying principles and mechanisms of LogProber, we first validate it through dedicated experiments in which we fine-tune a LLM (LLaMA; Touvron et al.
+(2023)) on a small set of cognitive question items taken from a recent study Yax et al.
 (2024a).
-Other methods have been developed that generally work on the assumption that contaminated text should be associated with highest completion probability.
-Some approaches focus on finding if a part of a document was included in the training corpus by looking if its token probability is particularly high Meeus et al.
-(2023) or by verifying if a permutation of the lines in the document is more likely than others Oren et al.
-(2024).
-However, this approach is ill-suited to evaluate contamination for some LLM benchmarks or more specifically cognitive tests, because they usually involve only a few very short questions as they require a lot of text to be efficient.
-Another very relevant approach, also based on token probability analysis is more precise and can be used with rather short texts Dong et al.
-(2024).
-It focuses on quantifying how many different completion a LLM is likely to generate to a given question, the lower the more likely the LLM is to be contaminated on the given question.
-However, we believe the fact that a LLM completes (or responds) always in the same way does not necessarily imply contamination, but could rather reflect higher confidence for a specific completion or response (which is often the case for answers provided in a question/answer format, such as the one often implemented in some benchmarks and often in cognitive tasks).
-The first aim of this paper is to present a new algorithm (LogProber) that allows for disentangling contamination from confidence in a way that is computationally cheap and adapted to the short question/answer format, such as the one usually used in machine learning benchmarks and cognitive tests.
-After illustrating the underlying principles and mechanisms behind LogProber, we validate it in dedicated experiments where we finetune a LLM (LLAMA Touvron et al.
-(2023b)) with specific items of a recently published cognitive test Yax et al.
-(2024).
-The results show that the method is effective in detecting contamination.
-However, in additional model training experiments, where the model doesn’t fit on all the question and answer tokens but only on the answer tokens, we found that LogProber was incapable of detecting the contamination.
-As such, the second aim of the paper is to demonstrate that not all types of contamination can be detected by simple analysis of token probability of a given benchmark or cognitive test question.
-Intriguingly, our fine-tuning experiments also show that LLMs can perfectly memorize the answer to a given question without showing signs of memorization of the question itself.
+The results show that our method is effective in detecting contamination when it involves both the question and its answer.
+However, in additional model training experiments, we also highlight the inherent limitations of the algorithm in cases where contamination involves only the answer.
+Finally, we extend our analysis with another set of dedicated experiments using a wide range of questions from standard benchmarks (MMLU Hendrycks et al.
+(2021) and BigCodeBench Zhuo et al.
+(2025)), compare LogProber’s performance to that of CDD, and conclude by discussing their relative advantages and weaknesses across different types of contamination.
 2 Methods
 2.1 What is contamination and how does it relate to confidence?
-Benchmarks and cognitive tasks are usually collections of questions (’Q’) associated with a specific correct answer (’A’).
-LLMs’ performance is usually assessed giving ’Q’ as input (or context, prompt) and taking the resulting completed text as the response of the model.
-The performance of the LLM is then assessed by comparing its response with the known correct answer (’A’).
-Contamination broadly refers to the fact that the testing materials (benchmark questions or cognitive tasks) were present in the training corpus.
-LLM’s performance is then considered to be driven directly by training data rather than from its generalization (or emergent cognitive) capacities.
-In the context of benchmarks and cognitive tasks, the question of contamination boils down to answering the question of whether or not the model has been trained on the test material (’Q’ and ’A’), because if the sample ’Q-A’ appears in the training dataset, the model will learn to "automatically" predict A from context ’Q’.
-In the first part of the paper, we assume that the model is trained of the full sequence ’Q-A’ as opposed to only finding the question ’Q’ in the training data for example without the associated answer.
-We will challenge this assumption later in the paper and discuss what happens if this is not the case.
-Contamination (i.e., the presence of a given ’Q-A’ in the training corpus), will make the generation of the response ’A’ very probable after the context ’Q’.
-In analogy with human cognitive science Efklides (2006), LLM’s response probability to a given query is considered to reflect the confidence of the model.
-It is therefore important to clarify here the difference and the relation between these key concepts of contamination and confidence.
-First, contamination is a state of the model (said to be contaminated) deriving from the relation between its training corpus and the testing material.
-Second, confidence is a feature of the model’s response to a given answer (said to be confident).
-Crucially, while it is true that responses ’A’ to contaminated ’Q-A’ pairs will generally appear very confident, higher confidence can only be achieved by other means.
-To understand this, consider the case where an LLM is asked the following:
-"What is the next character in the sequence 17817817817817817?".
-Many models will complete with "8" with a very high probability (i.e., confidence). However, it is very unlikely that this particular ’Q-A’ was present in the training corpus.
+Benchmarks typically consist of collections of questions (’Q’) paired with correct answers (’A’).
+LLM performance is usually assessed by providing ’Q’ as input (e.g., as context or a prompt) and evaluating the model’s generated response.
+This response is then compared to the known correct answer (’A’) to determine the model’s accuracy.
+Contamination broadly refers to the presence of testing materials (benchmark questions or cognitive task items) in the training corpus.
+In such cases, the LLM’s performance may be attributed directly to memorization of the training data, rather than to genuine generalization or emergent cognitive capabilities.
+In the context of benchmarks and cognitive tasks, the key question becomes whether the model was trained on the test material—specifically, on the ’Q-A’ pairs.
+If a full ’Q-A’ sample appears in the training data, the model may learn to predict ’A’ from ’Q’ automatically, without demonstrating real "understanding" (or rather generalization).
+In the first part of this paper, we assume that the model has been trained on the complete ’Q-A’ sequence, as opposed to having encountered only the question ’Q’ or the answer ’A’ in isolation.
+Later in the paper, we will challenge this assumption, exploring other forms of contamination.
+Contamination (i.e., the presence of a given ’Q-A’ pair in the training corpus) makes the generation of the response ’A’ following the context ’Q’ highly probable (this is the logic behind the CDD method).
+However, contamination is not the only reason a LLM might produce a highly probable response.
+In fact, by analogy with human cognitive science, a LLM’s response probability to a given query is generally interpreted as a reflection of the model’s confidence Efklides (2006).
+It is therefore important to clarify the distinction and relationship between the key concepts of contamination and confidence.
+Contamination refers to a property of the model — said to be contaminated — arising from a leakage of the testing material in its training corpus.
+Confidence refers to a property of the model’s response to a given input — namely, that the model’s answer is confident.
+Crucially, while responses to contaminated ’Q-A’ pairs will generally appear highly confident, high confidence can also be achieved in other ways.
+To illustrate this, consider prompting a LLM with the following question:
+"What is the next character in the sequence 17817817817817817?"
+Many models will complete the sequence with "8" with very high probability (i.e., confidence), due to their well-demonstrated inferential and in-context learning abilities Brown et al.
+(2020a).
+However, it is highly unlikely that this specific ’Q-A’ pair was present in the training corpus.
 2.2 A way to measure contamination in LLMs: the LogProber algorithm
-Most models are accessed in a ’black box’ setting, where we do not have access to the training corpus nor the model’s weights, we can only infer contamination from the probabilities of the generated tokens.
+Going back to our goal of finding contamination in LLMs, most models are accessed in a ’black box’ setting, where we do not have access to the training corpus nor the model’s weights; we can only infer contamination from the probabilities of the generated tokens.
 If the sequence ’Q-A’ appears in the training data, the model is contaminated by it, thus the probability of responding ’A’ from ’Q’ should be high.
-Such probability could be estimated either by directly accessing the tokens’ log(probabilities) or by running multiple queries and deriving "empirical" distribution of ’A’.
+Such probability could be estimated either by directly accessing the tokens’ logprobabilities or by running multiple queries and deriving "empirical" distribution of ’A’.
 Then, as proposed in Dong et al.
 (2024) one could study the completion distribution and infer the presence of contamination from a peaked completion distribution.
-However, this approach does not rule out that an LLM responds ’A’ to a given ’Q’ with high confidence because of its inferential and generalization capabilities.
-Indeed the ’A’ tokens probabilities mix confidence and contamination behaviors making the measure unreliable to disentangle between the two.
+However, this approach does not rule out that a LLM responds ’A’ to a given ’Q’ with high confidence because of its inferential and generalization capabilities.
+Indeed the ’A’ token probabilities mix confidence and contamination behaviours making the measure unreliable to disentangle between the two.
 To tackle this issue, our approach proposes an alternative method that switches the focus on ’Q’ instead of ’A’.
-Indeed while it is possible to find a peak in the generation distribution of ’A’ due to a high confidence (and not because of contamination), it is more unlikely that the model is able to predict the question ’Q’ confidently without having been explicitely trained on it.
+Indeed, while it is possible to find a peak in the generation distribution of ’A’ due to a high confidence (and not because of contamination), it is more unlikely that the model is able to predict the question ’Q’ confidently without having been explicitly trained on it.
 Assuming we can access the probability of each consecutive token in a given ’Q’ sequence, we can plot the cumulative log-probability of this sequence.
 This graph can be understood as the ’surprise’ the model has when generating a new token after all the previous ones.
-Crucially, for ’Q’ present in the corpus we can expect the probability of the subsequent tokens to approach p=1 (signalling that the token is predicted with high confidence) and therefore the log-probability will quickly achieve a plateau.
+Crucially, for ’Q’ present in the corpus we can expect the probability of the subsequent tokens to approach p=1 as soon as the LLM recognizes the sentence and therefore the cumulative log-probability will quickly achieve a plateau.
 To understand this point see for example what the shape of a very frequent phrase (e.g., the first line of a very famous and old poem Divine Comedy " When half way through the journey of our life"; Figure 1) should look like compared to a (semantically equivalent), but not as frequent, sentence ("Once at the middle of the voyage of my life" Figure 1; right).
-In the second case, successive tokens will be comparatively more surprising (or predicted with less confidence) (see 1).
-Once obtained the log-probability curves of ’Q’ sequences, their shape can be analysed in order to quantify how much they plateau to get insight into whether they were the model was ’Q’ was present in the corpus or not.
-More specifically, we propose to fit a very simple 2 parameters function to curve:
-with 2 variables (asymptote) and (acceleration).
-If the model is surprised by each new individual token in the sequence, it should result in a linear graph that will make the model slightly diverge by returning very high A and low B.
-Conversely, if the model predicts each consecutive token with high probability resulting curve "saturates" very quickly (low A and high B), therefore meaning that is likely that the ’Q’ sequence was present in the training corpus (contamination).
+In the second case, successive tokens will be comparatively more surprising (see 1).
+Once the log-probability curves of ’Q’ sequences are obtained, we can analyze their shape in order to quantify how much they plateau to get insight into whether ’Q’ was present in the corpus used to train the model.
+To measure this plateau in the cumulative logprobability graph, we propose to sort these cumulative logprobability values in increasing order.
+In this way, if the graph plateaus a lot, the first part of this list should have very low values, and only the very end should have high values, leading to a very small area under the curve.
+On the other hand, if the graph does not plateau and keeps increasing steadily, the area under the curve should be much larger.
+As such, the algorithm we propose here consists of sorting the cumulative logprobabilities in increasing order and estimating the area under the curve by summing the sorted values (step function approximation of the integral).
+In practice, this integral should have a lot of variance between contaminated and uncontaminated sequences.
+Thus, we will work with the logarithm of this integral instead that we call Safe Score.
+Experimentally we found that if the Safe Score is lower than 1, the item is likely contaminated.
+The algorithm is further detailed in Appendix C and is illustrated in Figure 1.
 2.3 Predictions and training experiment
-By returning the A and B parameters, LogProber quantifies the shape of the log-probability curve for each sentence and quantitative insights related to how much the model is familiar with the given question.
-These parameters can be very high or very low depending on the graph they are fitted on.
-As such we will consider their logarithm to reduce the variance : log(A) and log(B).
-This quantity is then used as a proxy for contamination on the answer (even if indeed having seen the question in the training corpus doesn’t mean the model has been exposed to its answer - we will discuss this later).
-To test the validity of this method and the LogProber algorithm, we designed several experiments.
-First, we compared the log(A) and log(B) parameters obtained by fitting ’Q’ in a cognitive test that has been published decades ago (the Cognitive Reflection Test; oldCRT; Toplak et al.
-(2013); Frederick (2005a)) to those of structurally and conceptually similar ’Q’ which have been designed after the creation of the model and cannot have been included in the training corpus (newCRT; Yax et al.
-(2024)) (see Table 1).
-We predicted that the log(A) and log(B) parameters of the oldCRT and the newCRT should be distinguishable (e.g., lay in different regions of the log(A) and log(B) plane).
+To test the validity of this method, we designed several experiments.
+We first show the algorithm works to spot contamination in LLMs and we then compare it with CDD Dong et al.
+(2024) on tasks where the model is susceptible to exhibit high confidence and then on more general topics.
+In the first experiment, we compared the Safe Score of LogProber obtained by fitting ’Q’ in a cognitive psychology test.
+Despite not being well known in the machine learning literature, these tests are particularly interesting when testing contamination as they are rather short, there are very few of them but each of them has been used very much in the literature and on forums so that most LLMs know them by heart.
+We decided to use one of the most famous : the Cognitive Reflection Test (noted as oldCRT; Toplak et al., 2013; Frederick, 2005).
+To compare the results obtained on these very likely known items, we used a version designed for LLMs published recently with reframed items that models are not familiar with (noted as newCRT; Yax et al., 2024a) (see Table 2 in appendix B).
+To be sure the LLM used is not familiar with these new items, we decided to focus on Llama 1 7B Touvron et al.
+(2023) that was trained well before the publication of these reframed items.
+Secondly, after testing LogProber in the wild, we test it in a controlled environment by contaminating a LLM on purpose on the newCRT items.
+Thirdly, we explore more fine-tuning scenarios by testing situations where the model is not trained on the ’Q-A’ sequence but only on the question (noted ’Q’ scenario) or only to produce the correct answer to a given question (noted ’A’ scenario).
+These experiments further allowed us to explore the complex relationship between contamination and the performance of a model.
+Finally, we compare LogProber with CDD Dong et al.
+(2024) first in a MCQ task with a top-end language model likely to be very confident on this task.
+Then we compare both in a less specific scenario.
+| Type of training | Description | Example |
+| STD-Training | is the original Alpaca training Taori et al.
+(2023) with no data from new CRT.
+We only simplified the instruction prompting (see Appendix A).
+| Q: A bat and a ball cost £1.10 in total.
+The bat costs £1.00 more than the ball.
+How much does the ball cost?
+A: The ball costs £0.05.
+|
+| QA-Training | involves the alpaca training with new CRT pairs of questions and answers.
+Thus the model is trained to predict each token of the question as well as each token of the answer given the question.
+| Q: A bat and a ball cost £1.10 in total.
+The bat costs £1.00 more than the ball.
+How much does the ball cost?
+A: The ball costs £0.05.
+|
+| A-Training | only fits on the answer given the question.
+Thus the model is not optimized the predict the question’s tokens but only the answer’s.
+| Q: A bat and a ball cost £1.10 in total.
+The bat costs £1.00 more than the ball.
+How much does the ball cost?
+A: The ball costs £0.05.
+|
+| Q-Training | only learns the question’s tokens and therefore is only trained to predict the question’s tokens and not the answer.
+| Q: A bat and a ball cost £1.10 in total.
+The bat costs £1.00 more than the ball.
+How much does the ball cost?
+A: The ball costs £0.05.
+|
+3 Results
+3.1 Can LogProber discriminate between newly designed and old ’Q-A’ items ?
+To validate LogProber, we first deployed it on "classical" items of the CRT (oldCRT) and to more recently developed ones (newCRT).
+NewCRT items have been published recently, thus they should not be represented in the training corpus of the considered LLM Yax et al.
+(2024a).
+The analysis shows that the Safe Score for the 7 CRT items remains in different parts of the score line (Figure 2 upper part).
+More specifically, the score was significantly lower (oldCRT: -0.40 newCRT: 2.64 t(10)=9.70, p<0.001) in the oldCRT compared to the newCRT test.
+To sum up, LogProber was able to differentiate oldCRT items (present in the training corpus) from newCRT items (absent in the training corpus).
+3.2 Can LogProber discriminate between newly designed ’Q-A’ items before and after contamination ?
+In the previous analysis, we showed promising evidence that LogProber is capable of discriminating between contaminated items and some conceptually equivalent, non-contaminated, versions (observational experiment).
+In the present section, we aim to provide experimental support to our claim by running dedicated fine-tuning experiments.
+To do so, we finetuned Llama 7B using both the questions and its answers (’Q-A’ tuning, see Table 2) on the newCRT items, which corresponds to contaminating the model with them.
+Results are reported in Figure 2 and show that the Safe Score of LogProber lies in very different sectors, before and after fine-tuning.
+The fine-tuning leads to low Safe Scores in the contaminated experiment (QA) as well as very high accuracies.
+On the other hand, in the control experiment, scores are very low and accuracy is average.
+This experiment validates the fact that scores capture model contamination if the model is contaminated in a Q-A manner.
+More specifically, the Safe Score was significantly lower (STD=2.98, QA=-0.83, t(8)=9.57, p<0.001) in the contaminated version (QA) compared to the original LLM test (STD).
+3.3 What is the simultaneous impact of different fine-tuning strategies on accuracy and contamination markers?
+Our first finetuning experiment featured a ’Q-A’ type of training, where the model was forced to learn to predict both the question and the answer of each individual CRT item.
+However, this represents only one way in which a model can be contaminated (See Table 2).
+For instance, the model can be trained on predicting the answer, given the question (without being trained to memorize the question itself).
+Such ’-A’ training strategy is very common to fine-tune the model on instructions for example Zheng et al.
+(2023); Taori et al.
+(2023).
+Conversely, a model could be trained only to predict the question, without necessarily including the (correct) answer in the string.
+Such ’Q-’ training strategy style would correspond to the presence in the corpus of a given question without the corresponding answer, for example when it is quoted to explain it or in research papers.
+In Figure 3 we report contamination scores for both these types of training as well as accuracy and found that, indeed, the contamination log-scores do not capture the contamination for the -A type of training as the model doesn’t memorize the question.
+As for the Q- style of training, we found that the contamination scores are very high while the model’s accuracy doesn’t increase with the training (the model is contaminated on the question and not the answers).
+This series of experiments shows the complex relationship between markers of contamination (as derived from LogProber algorithm) and accuracy.
+On one side, (and somehow unsurprisingly) the accuracy is high whenever the model is trained to predict ’A’ given ’Q’, regardless of whether or not the ’Q’ itself is contained in the training.
+A language model can know by heart the answer to a question without knowing the question itself and can even show signs of surprise when reading the question.
+On the other side, LogProber detects contamination whenever the model has been trained to predict ’Q’, regardless of whether or not the ’A’ was included.
+3.4 Comparison with CDD
+As we believe contamination can be difficult to disentangle from confidence only looking at the answer, LogProber switches focus to detecting familiarity with the question rather than on the answer only.
+While this design works very well on QA types of finetuning, we have just shown its limitations on Q- and -A types of finetuning.
+It is now natural to compare this method with CDD, that only uses the answer to see how these two paradigms of contamination detection behave in practice on classical machine learning benchmarks.
+As such, we first used MMLU Hendrycks et al.
+(2021) a MCQ type benchmark where we contaminated Qwen2.5-32B-Instruct Qwen et al.
+(2025) on part of the test set either following a QA (contaminated) or STD (control) type of training using QLoRA Dettmers et al.
+(2023).
+Our prediction is that Qwen2.5, being a top-end model, should be very good at MMLU thus should show very high confidence on this task.
+We separated a part of the test set into two splits : split A is included in the training data, while split B is not.
+Results are shown in Figure 4 left side.
+As expected, while LogProber is able to differentiate split A from split B in QA, CDD considers all splits to be contaminated on MMLU in all finetuning conditions even if the model is not contaminated (STD condition).
+See Appendix D for detailed comparison metrics, training details and different learning rates.
+Now let’s compare both algorithms in a different task where models should show less confidence in answering the question.
+We contaminated a model on part of the BigCodeBench dataset, which contains coding problems where LLMs need to write a function corresponding to the given instructions.
+Figure 4 right side shows the results on the coding task.
+In this situation, CDD performs much better than in MMLU but seems still less effective than LogProber.
+See Appendix D for the detailed comparison metrics.
+Even if LogProber appears on top of CDD in these experiments, it is important to emphasize the fact that here we only limited ourselves to the QA finetuning scenario.
+CDD would greatly outperform LogProber on Q- or -A types of finetuning.
+4 Discussion
+In this paper, we propose and validate a simple algorithm, LogProber, which is capable of detecting common forms of contamination in LLMs while disentangling confidence from contamination.
+It does so by analyzing the log-probability of the question, rather than focusing on the answer.
+This makes LogProber highly efficient compared to other approaches, as it requires only a single forward pass on the question.
+By comparison, CDD Dong et al.
+(2024) requires multiple answer generations—although it is worth noting that CDD does not require access to log-probabilities.
+We demonstrated the effectiveness of LogProber using both observational and experimental evidence.
+For the observational evidence, we show that LogProber can distinguish between the widely known items of a popular cognitive experiment—the Cognitive Reflection Test—and newly designed variants of those same items Frederick (2005); Toplak et al.
+(2013); Yax et al.
+(2024a) .
+More specifically, the contamination scores for the new items were significantly different from those of the original items.
+As for the experimental evidence, we conducted dedicated experiments in which we fine-tuned a model on the ’Q-A’ sequences of newCRT items.
+These experiments demonstrated that LogProber was able to successfully detect contamination in the fine-tuned model.
+We also conducted additional experiments involving different fine-tuning scenarios: training the model to predict only the question (‘Q-’ training) or only the answer (‘-A’ training).
+Critically—and somewhat unsurprisingly—our results showed that LogProber was able to detect contamination only when the question was included in the training data.
+Third, we evaluated the method on classical machine learning benchmarks (MMLU and BigCodeBench), demonstrating near-perfect contamination detection in our experiments, with accuracy exceeding 98% in our experiments.
+In the same setting, CDD was less effective, frequently misidentifying uncontaminated scenarios as contaminated.
+This is due to the fact that CDD relies on the probability of the generated response, which can inherently conflate response confidence with contamination.
+LogProber avoids this issue by evaluating the probability on the question, rather than focusing solely on the answer.
+However, it is also important to note that question-based methods like LogProber are well-suited for detecting ’Q-A’ contamination—typically introduced during pretraining—but are less effective at identifying answer-only (’-A’) contamination, which is more likely to occur during fine-tuning Zheng et al.
+(2023).
+To sum up, question-based methods (such as LogProber) and answer-based methods (such as CDD) offer complementary advantages and limitations.
+Question-based methods can detect contamination originating during pretraining, providing valuable insights into the contents of the training corpus.
+However, they are generally unable to detect contamination involving only the answer, which typically arises during fine-tuning.
+Conversely, answer-based techniques struggle to disentangle contamination from confidence.
+As a result, they are less reliable for models that are genuinely strong or highly confident—an increasingly common situation with top-tier models on classical benchmarks like MMLU.
+An interesting direction for future work is to combine these two approaches, aiming to leverage the strengths of each.
+However, this is not as straightforward as it may seem.
+For an answer-based method to flag contamination, the model must either be trained on the full ’Q-A’ pair, trained on the answer alone (’-A’ training), or simply be confident in the answer.
+For a question-based method to flag contamination, the model must have been trained on either the full ’Q-A’ pair or the question alone (’Q-’ training).
+As a result of these limitations, when evaluating the performance of a LLM, it is not possible to attribute its accuracy with absolute certainty to either high confidence or contamination.
+However, different combinations of outcomes from the question-based and answer-based methods can help constrain our interpretation of the origins of the LLM’s performance.
+If both methods return positive (i.e., suggest contamination), the model may have been trained on the full ’Q-A’ pair, or it may have seen the question (’Q-’ training) and happens to be confident in its answer.
+(We cannot be 100% certain without additional assumptions).
+If the question-based method is positive but the answer-based method is negative, it suggests the model was trained on the question alone.
+If the answer-based method is positive but the question-based method is negative, it likely indicates ’-A’ training—i.e., the model was exposed to the answer during fine-tuning without seeing the question or that the model is just confident in its answer.
+Finally, if both methods return negative, we can reasonably conclude that the model was not contaminated on the item.
+In summary, while answer-based methods cannot rule out the possibility that a model is simply confident, and question-based methods cannot rule out the possibility that a model has been fine-tuned in a specific manner, combining both approaches allows for more educated guesses about the presence and type of contamination and the origin of its performance.
+To conclude, this work makes three main contributions.
+First, we propose and validate a method for estimating the likelihood of contamination in short text sequences, based on a computationally efficient analysis of sequence log-probabilities.
+Our approach, LogProber, is particularly well-suited for detecting contamination likely to occur during pretraining (’Q-A’ training).
+Second, we clarify an important distinction between contamination and confidence, and show how different types of contamination (’Q-A’, ’Q-’, and ’-A’) interact with different detection methods (question-based and answer-based), each with varying effectiveness depending on the fine-tuning scenario.
+Third, we discuss the upsides and drawbacks of question-based and answer-based approaches of contamination detection methods and how taking into account training techniques and model behavior is vital to account for contamination detection.
+This work could pave the way for new ways of detecting and thinking about contamination in efficient and effective manners—something much needed given the ever faster-evolving landscape of LLMs and the resulting need to monitor the evolution of their performance Yax et al.
+(2024b).
+5 Limitations
+A limitation that we did not discuss yet of our approach is that it requires having access to the log-probabilities for the tokens and the question of the Q-A pair (typically used in benchmark and cognitive tasks).
+It could be, in principle, extended to cases where this information is not present by sampling each token in a sequence and deriving an "empirical" distribution.
+We note however that in that way the method would become more computationally expensive.
+We would also like to emphasize the fact that LogProber focuses on how easily the LLM can predict a question, which is common when the model is contaminated.
+However, LLMs might also predict questions through in-context learning.
+For example, take the 178178… question — while the start is hard to predict, once the pattern emerges, a LLM can predict the end, resulting in a flatter line on the graph.
+This can be an issue for sequences with strong patterns, but in practice, such questions still score higher in Safe Score than truly contaminated ones, as the graph plateaus more slowly.
+Finally, LogProber presents other limitations, for instance, the algorithm returns a score that can be difficult to interpret per se.
+To be interpreted, they have to be compared with other sequences to put it in perspective and decide how likely the model is contaminated on the sequence.
+In our paper, we set the threshold between contamination and Safe to 1 as it seemed to work very well in almost all our experiments.
+We also did not test LogProber in implicit contamination scenarios, where the LLM is trained on questions very close to the one tested.
+This could be done in future work in order to test question-focused contamination detection techniques on this type of contamination and compare the results with algorithms like CDD that perform even on implicit contamination only looking at the answers.
+Lastly, while we took care of using a model that was not familiar with our items in our first two experiments with Llama 1 7B, in the 3 experiment we used Qwen2.5-32B-Instruct that was developed after the publication of MMLU.
+We thus cannot be sure the test set we used to measure the contamination of the model already contained sequences the model was familiar with.
+We can at least claim that they were probably not in the pretraining set as LogProber didn’t catch any of these as contaminated.
+However, CDD might have caught some of them if indeed the model was trained on the MMLU test set during -A style finetuning during its post-training.
+6 Acknowledgements
+This work was granted access to the HPC/IA resources of [IDRIS HPE Jean Zay A100] under the allocation 2024- [AD011013693R2] and [IDRIS HPE Jean Zay H100] under the allocation 2024-[A0171011996] and made by GENCI.
+SP is supported by the European Research Council under the European Union’s Horizon 2020 research and innovation program (ERC) (RaReMem: 101043804), the Agence National de la Recherche (CogFinAgent: ANR-21-CE23-0002-02; RELATIVE: ANR-21-CE37-0008-01; RANGE: ANR-21-CE28-0024-01; ANR-17-EURE-0017), the Alexander Von Humboldt foundation, the EUR Frontiers in Cognition, the Idex PSL and a Google unrestricted gift.
+PYO is supported by ANR AI individual chair ANR-19-CHIA-0004.
+Appendix A Llama 1 7B finetuning details
+The classic Alpaca finetuning procedure is composed of 52k examples obtained from text-davinci-003 using the self-instruct pipeline Taori et al.
+(2023).
+They are composed of 3 fields : Instruction, Input and Response prompted with the following format :
+Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
+### Instruction:
+{instruction}### Input:
+{input}### Response:
+{response}
+The model only fits on tokens from the Response field. The input field is omitted if no input is provided for the given instruction.
+To contaminate the model on newCRT we simplified the prompting to :
+{instruction}
+{input}
+{response}
+and included both the question and the answer in the response field so that the model fits on both of them.
+The point of this procedure is to make the model fit on the newCRT tokens from scratch without prompting coming before as we don’t want the model to condition the generation of newCRT items to the alpaca instruction prompting scheme.
+While the alpaca default training data only fit on the response we formatted the newCRT data in different manner for the 3 training methods.
+For the QA-Training we included both question and answer in the response field so that the model fits on both.
+For the A-Training we included the question in the instruction field and the answer in the response field.
+Finally for the Q-Training we only included the question in the response field without the answer so that the model only fits on the question tokens.
+All these finetuning were done by finetuning all the weights of the model.
+Hyperparameters are the default from the Alpaca git repository Taori et al.
+(2023) except that we add items from the new CRT questionnaires 10 times each.
+The questionnaire contains 7 questions and the model is trained for 3 epochs.
+Therefore, the model will see each of the 7 question-answer pair 30 times among the 52 000 Alpaca examples during the training process.
+These were run on a single GPU A100 taking 5 to 8 hours depending on the training set for each run.
+Appendix B Old CRT and new CRT questionnaires
+The list of items referred to as old CRT and new CRT are borrowed from Yax et al.
+(2024a) where the authors used original items from Frederick (2005) which language models are likely to be familiar with and made new fresh ones after the training of these models (thus LLMs trained before that time cannot be contaminated on them).
+Items are shown in Tab 2.
 | oldCRT items | newCRT items |
@@
 |
-Second, we compared log(A) and log(B) parameters of the newCRT items in a "native" LLM (i.e., uncontaminated by the ’Q-A’ pairs; llama 1 7B model Touvron et al.
-(2023b)) with a version of the same model fine-tuned and contaminated (using the Alpaca pipeline Taori et al.
-(2023)).
-Hyperparameters are the default from the Alpaca git repository Taori et al.
-(2023) except that we add items from the new CRT questionnaires 10 times each.
-The questionnaire contains 7 questions and the model is trained for 3 epochs meaning the model will see each of the 7 question-answer pair 30 times during the training process.
-In addition, we simplified the prompting and we saved the models after each epoch (see Appendix A for more details).
-We predicted that the log(A) and log(B) parameters of the newCRT items should change after fine-tuning, more specifically, as the fine-tuned model becomes contaminated with ’Q-A’ the log(A) parameter should decrease and the log(B) parameter should increase (QA-training).
-Type of training | Description | Example |
-STD-Training | is the classic Alpaca training with no additional data from the new CRT.
-We replicated similar training conditions than in the original Alpaca paper Taori et al.
-(2023), no additional data but with our simplified instruction prompting.
-As explained above, this model hasn’t be exposed to the newly reframed CRT items therefore it shouldn’t be contaminated.
-| Q: A bat and a ball cost £1.10 in total.
-The bat costs £1.00 more than the ball.
-How much does the ball cost?
-A: The ball costs £0.05.
-|
-QA-Training | involves fitting on both the questions and the answer pairs of the new CRT questionnaire.
-This means the model is trained to predict each token of the question given the previous tokens as well as each token of the answer given the question and the previous tokens in the answer.
-| Q: A bat and a ball cost £1.10 in total.
-The bat costs £1.00 more than the ball.
-How much does the ball cost?
-A: The ball costs £0.05.
-|
-A-Training | only fits on the answer given the question.
-Thus the model is not optimized the predict the question’s tokens but only the answer’s.
-| Q: A bat and a ball cost £1.10 in total.
-The bat costs £1.00 more than the ball.
-How much does the ball cost?
-A: The ball costs £0.05.
-|
-Q-Training | only fits on the question’s tokens and therefore is only trained to predict the question’s tokens and not the answer.
-| Q: A bat and a ball cost £1.10 in total.
-The bat costs £1.00 more than the ball.
-How much does the ball cost?
-A: The ball costs £0.05.
-|
-Finally, we run two additional fine-tuning experiments, featuring hybrid fine-tuning scenarios in which, for each item of the cognitive test, the model was only fine-tuned in the question ’Q’ (Q- training) or the ’A’ (A- training) (see Table 2).
-These experiments further allowed us to explore the complex relationship between contamination and performance of a model.
-3 Results
-3.1 Can LogProber discriminate between newly designed and old ’Q-A’ items ?
-The Cognitive Reflection Text (CRT) Frederick (2005b) is one of most popular psychological questionnaires used to investigate human reasoning and typically associated to the demonstration of two modes of reasoning, the fast "system 2" and slow "system 1" Kahneman (2011).
-The test is widely used in the scientific literature and can be found even beyond in numerous journal articles, blogs and forums.
-Therefore it is very likely that LLMs trained on web data containing the CRT.
-To validate LogProber, we deployed it on "classical" items of the CRT (oldCRT) and to more recently developed ones (newCRT), which has been published only once and so recently so they should not be represented in the training corpus of the considered LLMs Yax et al.
-(2024).
-The analysis shows that the A and B parameters for the 7 CRT items lied in linearly separable sectors of the log(A)/log(B) plane (Figure 2).
-More specifically log(A) was significantly higher (oldCRT: -1.04 newCRT: 1.28 t(11)=6.29, p<0.001) and log(B) was significantly lower (oldCRT: 3.38 newCRT: 0.37 t(9)=5.83, p<0.001) in the oldCRT compared to the newCRT test.
-Looking more in details, LogProber returns higher scores for the first 3 oldCRT items, which are the first having being published and more wide-spread in the literature Frederick (2005a); Toplak et al.
-(2013).
-To sum up, LogProber was able to differentiate oldCRT items (present in the training corpus) from newCRT items (absent in the training corpus).
-These first results therefore provide first observational evidence of the effectiveness of LogProber.
-3.2 Can LogProber discriminate between newly designed ’Q-A’ items before and after contamination ?
-In the previous analysis, we showed promising evidence that LogProber is capable of discriminating between contaminated items and some conceptually equivalent, non-contaminated, versions (observational experiment).
-In the present section, we aim to provide experimental support to our claim by running dedicated fine-tuning experiments.
-To do so, we finetuned Llama 7B using both the questions and its answers (’Q-A’ tuning, see Table 2) the newCRT items, which corresponds to contaminating the model with them.
-Results are reported in Figure 3 and show that the log(A) and log(B) parameter of LogProber lie in very different sector of the the plane, before and after fine-tuning.
-The fine-tuning lead to high contamination log scores in the contaminated experiment (QA) as well as very high accuracies.
-On the other hand, in the reference experiment scores are very low and accuracy is average.
-This experiment validates the fact that contamination scores capture heavy model contamination if the model is contaminated in a Q-A manner.
-More specifically log(A) was significantly higher (STD=2.09, QA=-1.38, t(9)=5.19, p<0.001) and log(B) was significantly lower (STD=-0.43, QA=3.50, t(11)=4.45, p<0.001) in the fine-tuned compared to the native LLM test.
-These results show that in a carefully controlled experiment, where we explicitely manipulated the training of a model to include the newCRT items (Q-A), LogProber was capable to detect that the model was finetuned with the newCRT Q-A items.
-3.3 What is the simultaneous impact of different fine-tuning strategies on accuracy and contamination markers?
-Our first finetuning experiment featured a ’Q-A’ type of training, where the model was forced to learn to predict both the question and the answer of each individual CRT item.
-However, this represents only one way in which a model can be contaminated (See Table 2).
-For instance, the model can be trained on predicting the answer, based on the question (without further training the question per se).
-Such ’-A’ training strategy is very common to fine-tune the model on instructions Zheng et al.
-(2023); Taori et al.
-(2023).
-Conversely, a model could be trained only to predict the question, without necessarily including the (correct) answer in the string.
-Such ’Q-’ training strategy style would correspond to the presence in the corpus of a given question without the corresponding answer, for example when it is quoted to explain it or in research papers.
-In Figure 4 we report contamination scores for both these types of training as well as accuracy and found that, indeed, the contamination log-scores do not capture the contamination for the -A type of training as the model doesn’t memorize the question.
-As for the Q- style of training, We found that the contamination scores are very high while the model’s accuracy doesn’t increase with the training (the model is contaminated on the question and not the answers).
-This series of experiments show the complex relationship between markers of contamination (as derived from LogProber algorithm) and accuracy.
-On one side, (and somehow unsurprisingly) the accuracy is high whenever the model is trained to predict ’A’ given ’Q’, regardless of whether or not the ’Q’ itself is contained in the training.
-A language model can know by heart the answer to a question without knowing the question itself and can even show signs of surprise when reading the question.
-On the other side, LogProber detects contamination whenever the model has been trained to predict ’Q’, regardless of whether or not the ’A’ was included.
-4 Discussion
-In this paper we propose and validate a very simple algorithm, LogProber, which is capable to spot common forms of contamination in LLMs.
-More specifically, LogProber can be applied to detect contamination of text in a quesiton-answer (Q-A) format, which is commonly used both in benchmarks used in machine learning, as well as in psychology tests Srivastava et al.
-(2023); Yax et al.
-(2024); Coda-Forno et al.
-(2024).
-LogProber analyses the (log)probability of the question in a Q-A sequence and is computationally cheap because, compared to other approaches, LogProber only requires a single forward pass on the question and storing the log-probabilities of tokens.
-As comparison, Dong et al.
-(2024) requires several generations of the answer (but doesn’t require access to the log-probabilies).
-We demonstrated the effectiveness of LogProber by leveraging both observational and experimental evidence.
-Concerning observational evidence, we demonstrated the capacity of LogProber to differentiate between the classic items of a popular cognitive experiment (the Cognitive Reflection Test; Frederick (2005a); Toplak et al.
-(2013)) and newly designed versions of the same items Yax et al.
-(2024).
-More specifically, the contamination scores of the new items (in the form of log-parameters of an exponential function fitted in the log probabilities) were significantly different from those of the old items, which were, in turn, much better predicted by the LLMs.
-Concerning experimental evidence, we ran dedicated experiments where we trained (fine-tuned) a model to predict ’Q-A’ sequences of the newly designed items of the Cognitive Reflection Test.
-These experiments showed that LogProber was able to detect the changes in LLM training (i.e., it was capable of detecting its contamination).
-For the sake of completeness, we ran other experiments where the additional only concerned either learning to predict only the question (’Q-’ training) or only the answer (’-A’ training).
-Critically, and somehow unsurprisingly, our results showed that LogProber was able to detect contaminations only when the Question was included in the training.
-This result suggests that our approach is very effective in detecting contamination at the level of pre-training, since it is likely to occur in the ’Q-A’ format, and not in fine-tuning (likely to occur in the ’-A’ format: answer only) Wang et al.
-(2022); Taori et al.
-(2023); Zheng et al.
-(2023).
-This limitation does not substantially reduce the utility of LogProber, since in the open access LLM field pretraining dataset are rarely shared or very large to search a query in them.
-LogProber makes it possible to account for pretraining contamination at a very cheap cost.
-When training a model using a full language modeling method - a common approach for pretraining autoregressive decoder-only LLM Wang et al. (2022) — the training data can follow three scenarios:
-1.
-The model is trained only on the question (Q) if the data does not include the answer.
-2.
-The model is trained on both the question and the answer (QA) if they appear consecutively in the data.
-3.
-Neither the question nor the answer is included in the training data (STD).
-In this context, the QA scenario indicates data contamination.
-A low contamination score implies that neither Q nor QA scenarios are present in the training data, suggesting the STD scenario.
-Therefore, if a query results in a low contamination score, it is likely safe for use.
-Conversely, a high contamination score indicates that the model has been exposed to the question, suggesting potential contamination and that the model might be unreliable for that query.
-In the open access community finetuning datasets are often shared and a lot smaller making it possible to know explicitly if the model was trained on the given query.
-In short, LogProber can efficiently account for contamination acquired during pretraining (assuming the model is trained in a full language modelling - which is nearly always the case for modern LLMs) but may fail to capture finetuning contamination.
-This issue is mitigated by the fact that open access LLM finetuning datasets are often shared and fairly small making it possible to known explicitly if the model was trained on the given query.
-Combining LogProber and Dong et al.
-(2024) might improve the results as it would make it possible to better differentiate A from STD and Q from QA but it still suffers from this disentangling between contamination and confidence.
-Incidentally, these additional experiments also illustrate a very important, but often overlooked, fact: LLMs can be (over) trained to respond to a question, making them very familiar with answering the question without necessarily showing signs of knowing the question itself.
-Indeed, in our experiment ’-A’ training resulted in no sign of contamination, while the accuracy on the questions was extremely high.
-This also shows that LLM cognition is very different from human cognition especially in terms of how these two systems learn.
-Indeed, we can choose tokens on which LLMs learn while it is not possible to constraint the human learning mechanisms on specific words only.
-LogProber presents other limitations, for instance, the algorithm returns a scores that can be difficult to interpret per se.
-To be interpreted they have to be compared with other sequences to put it in perspective and decide how likely the model is contaminated on the sequence.
-In our paper, we focus on an unambiguous case, where newly designed were compared to structurally equivalent old CRT, which served as an obvious reference point.
-This method (designing new items) could not be easily scalable to large benchmarks.
-However, by comparing LogProber scores across similar benchmarks, it will nonetheless be possible to make hypotheses about the relative likelihood of contamination.
-Furthermore, future experiments featuring more questions and LLMs could provide in future insights concerning the mapping between the LogProber scores and the likelihood of contamination.
-This work also puts into perspective inherent limitations of quantifying contamination in LLMs in a black box setting on a given short sequence.
-Focusing only on the answer leads to mixing confidence and contamination while using the question solve this issue but only works in specific training settings.
-Further research could focus on doing both at the same time in order to increase the probability to get it right but we believe these methods would still not be foolproof and might require additional information such as neuronal activations from the network to accurately classify contaminated items.
-Another limitation of our approach is that it requires having access to the log-probabilities for the tokens and the question of the Q-A pair (typically used in benchmark and cognitive tasks).
-It could be, in principle, extended to cases where this information is not present by sampling each token in a sequence and deriving an "empirical" distribution.
-We note however that in that way the method would become more computationally expensive.
-In conclusion, we propose and validate a method to estimate the likelihood of contamination for speficic string of text, based on a very computationally cheap analysis of the log-probabilities of the sequence.
-Our approach, LogProber, is particularly suited to detect contamination that should occur at the level of pretraining and can be easily applied in a domain-general manner, as soon as the cognitive test of the benchmarks quetions are framed in the ’Q-A’ manner.
-Appendix A Training details
-The classic Alpaca finetuning procedure is composed of 52k examples obtained from text-davinci-003 using the self-instruct pipeline Taori et al.
-(2023).
-They are composed of 3 fields : Instruction, Input and Response prompted with the following format :
-Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request.
-### Instruction:
-{instruction}### Input:
-{input}### Response:
-{response}
-The model only fits on tokens from the Response field. The input field is omitted if no input is provided for the given instruction.
-To contaminate the model on newCRT we simplified the prompting to :
-{instruction}
-{input}
-{response}
-and included both the question and the answer in the response field so that the model fits on both of them.
-The point of this procedure is to make the model fit on the newCRT tokens from scratch without prompting coming before as we don’t want the model to condition the generation of newCRT items to the alpaca instruction prompting scheme.
-While the alpaca default training data only fit on the response (like the paper) we formatted the newCRT data in different manner for the 3 training methods.
-For the QA-Training we included both question and answer in the response field so that the model fits on both.
-For the A-Training we included the question in the instruction field and the answer in the response field.
-Finally for the Q-Training we only included the question in the response field without the answer so that the model only fits on the question tokens.
-Hyperparameters are the default hyperparameters for Llama 7B taken from the stanford alpaca git repository with a learning rate of 2e-5 and 3 epochs.
+Appendix C LogProber algorithm pseudocode
+The pseudo code for LogProber is provided in Algorithm 1.
+Appendix D Qwen2.5-7B-Instruct finetuning details and comparison with CDD
+To contaminate Qwen2.5-32B-Instruct on MMLU we selected 10,000 samples from the train set of cais/mmlu on the huggingface hub as well as 100 samples from the test set that we duplicated 100 times (split A).
+The split B is made of 100 other samples from the test set that are not included in the training set.
+Finetuning is performed using QLoRA Dettmers et al.
+(2023) instead of full finetuning, as we did in the other experiments, because this LLM is a lot larger, and this finetuning was performed using 1 H100 and lasted approximately 8 hours for each model.
+The pipeline and hyperparameters are the default from alpaca Taori et al.
+(2023) explained in Appendix A with 5 epochs instead of 3.
+We tested different learning rates.
+The main paper shows results for a learning rate of 5e-5.
+Figure 5 shows results for a learning rate of 5e-4 yielding very similar results (see Table 3 for metrics), and Figure 6 for a learning rate of 5e-6.
+In this last figure, the learning rate is very small which doesn’t change very much the way the LLM is generating the question text.
+As such, LogProber has a hard time noticing contamination (as it is very light on the question).
+However, CDD seems to maintain its accuracy on the coding benchmark.
+While this is an unexpected behavior we believe it could come from the fact that when predicting the question, for the first few tokens the LLM doesn’t have enough context to predict them accurately, this naturally increases the Safe Score of LogProber as the cumulative logprobability curse is less horizontal.
+However, the answer being generated after the question, the LLM has enough context to notice it knows the generation by heart which may lead it to retrieve answers more easily than questions.
+| Algorithm | Dataset | Learning Rate | Accuracy | Precision | Recall | F1 Score |
+| CDD | MMLU | 5e-4 | 0.50 | 0.50 | 1.0 | 0.67 |
+| LogProber | MMLU | 5e-4 | 1.0 | 1.0 | 1.0 | 1.0 |
+| CDD | MMLU | 5e-5 | 0.50 | 0.50 | 1.0 | 0.67 |
+| LogProber | MMLU | 5e-5 | 0.98 | 1.0 | 0.95 | 0.97 |
+| CDD | MMLU | 5e-6 | 0.5 | 0.5 | 1.0 | 0.67 |
+| LogProber | MMLU | 5e-6 | 0.57 | 1.0 | 0.13 | 0.23 |
+| CDD | CODE | 5e-4 | 0.63 | 0.58 | 0.95 | 0.72 |
+| LogProber | CODE | 5e-4 | 1.0 | 1.0 | 1.0 | 1.0 |
+| CDD | CODE | 5e-5 | 0.68 | 0.62 | 0.93 | 0.74 |
+| LogProber | CODE | 5e-5 | 0.99 | 1.0 | 0.99 | 0.99 |
+| CDD | CODE | 5e-6 | 0.76 | 0.68 | 0.97 | 0.80 |
+| LogProber | CODE | 5e-6 | 0.5 | 1.0 | 0.01 | 0.02 |
+A similar setup was made for the coding benchmark with the first file of MBXP Athiwaratkun et al.
+(2023) being used to extract 10,000 samples for the training set and then 100 samples were taken from the bigcode/bigcodebench dataset for the test set.
+Similarly, they were duplicated 100 times each for the training set for split A, and another 100 were used to compose split B (which are not added to the training data).
+For the STD training in MMLU we did only include the training set without splits A nor B making 10,000 datapoints in total.
+These were extracted from the question column for MMLU and we appended the answer column right after.
+For prompting reasons, because the question finishes with the "Answer:" prompt we added a left parenthesis afterwards to force the LLM to complete with the answer right after without reasoning.
+As the answer column starts with a left parenthesis we removed it to avoid having two of them.
+For the QA finetuning we included split A making 20,000 datapoints in total.
+For MBXP we used the content column for the train set and cut completions to the 1000th token in case they exceeded this length.
+For the splits A and B we used the complete_problem and canonical_solution columns appended after each other.
+In order to evaluate CDD we need to stop the generation at some point to define the limit of what is an answer to a given question.
+While some LLMs generate an end of text token by themselves it is not always the case and we noticed finetuned Qwen models did not always put one and started hallucinating after the answer.
+Thus, we added a prompt asking the model to stop after the answer, and, because it was not very effective in MMLU, we cut the answers to one character for this benchmark (the model generates the answer in the first token of the completion).
+For BigCodeBench we did not restrict the answers and let the model generate up to 100 new tokens.
+CDD hyperparameters were taken from the original paper Dong et al.
+(2024): is 0.05, is 0.01 and 50 answers were generated for each question with a temperature of 1 in addition to the greedily decoded answer.